OpenStudy (anonymous):

x3 = 1 over 8. 1 over 2 ±1 over 2 1 over 4 ±1 over 4

OpenStudy (mathmale):

It's important that you express exponentiation properly. In theory, your 'x3' means 3*x. If that '3' is an exponent, write the expression as x^3 = 1/8, or \[x^3=\frac{ 1 }{ 8 }~or~\frac{ 1 }{ 2^3 }\] and then solve for x.

OpenStudy (anonymous):

k thx

OpenStudy (mathmale):

Note that\[x^3=(\frac{ 1 }{ 2 })^3.\] What's the value of x?